# On the spectrum of Sturmian Hamiltonians of bounded type in a small coupling regime

### v1

@@ 3.5.2 Rectangles of Controlled Size
We will now use this base set of rectangles to construct dynamical rectangles of small controlled size. We will need rectangles of smaller scale to ensure that the Cantor set of $\mathcal{C}(\lambda,O_{r(\rho)})$, that we construct later on, will have small gaps in comparison to their bridges.
@@
    $\text{length}(\kappa)\leq C_{1}(\mu_{1}-\delta)^{-N}C^{*}$
@@
    $\text{length}(\kappa)\leq C_{1}(\mu_{1}-\delta)^{-N}\frac{C^{*}}{2}$
@@
    The non-stable boundaries $\partial^{ns}R^{\prime}_{m}(\lambda,Q_{i},N)$ of $R^{\prime}_{m}(\lambda,Q_{i},N)$ are line segments with endpoints in $\partial\tilde{K}(Q_{i})$. That is, $R^{\prime}_{m}(\lambda,Q_{i},N)$ has vertices belonging to $\partial\tilde{K}(Q_{i})$, and does not intersect $\tilde{K}(Q_{i})$ otherwise.
@@
Notice from (3.13) and Remark 2, provided $\lambda$ is small, the image of the cone $\tilde{K}(Q_{i})$, under the map $f_{\lambda,0,m}$, overflows the box $\mathbb{B}_{j}$ (see Figure 8). Without loss of generality, let us assume that this holds for all $\lambda\in[0,\lambda_{4}].$
@@ 3.6 Gap Estimates and Thickness
In order to establish (3.11), we will show that for each $\lambda\in[0,\lambda_{4}]$, there is a Cantor set $\mathcal{C}\left(\lambda,N^{*}\right)$ contained in $\mathcal{C}\left(\lambda,O_{r(\rho)}\right)$ that has thickness larger than $M$.
@@ 3.6.1 Gap Constructions
We will construct a Cantor set $\mathcal{C}\left(\lambda,N^{*}\right)$ that is contained in $\mathcal{C}\left({\lambda},O_{r(\rho)}\right)$ and show that it has thickness bigger than $M$:
@@
Let $J$ be such a component, and suppose it is contained in $R^{\prime}_{k_{0}+1}(\lambda,Q_{j},N^{*})$ for some $j=1,\dots,4$, and $J\cap O_{r_{0}}(Q_{j})\neq\emptyset$. Then, $J$ is the image of some segment $I_{0}^{\prime}\subset I_{0}$ under the map $f_{\lambda,k_{0},0}$. We will say that $\text{int}I^{\prime}_{0}$ is a gap of order $k_{0}$, where the interior is taken with respect to the subspace topology on $\hat{L}_{\lambda}$.
@@
$f_{\lambda,k_{0},0}\left(I_{0}^{\prime}\right)$ is a curve tangent to $\mathcal{K}^{u}$, by Remark 5 and our discussion after Lemma 12, as $f_{\lambda,k_{0},0}\left(I_{0}^{\prime}\right)\cap O_{r_{0}}(Q_{j})\neq\emptyset$, we must have that $f_{\lambda,k_{0},0}\left(I_{0}^{\prime}\right)$ is fully contained in $R^{\prime}_{k_{0}+1}(\lambda,Q_{j},N^{*})$ and only intersects $\partial R^{\prime}_{k_{0}+1}(\lambda,Q_{j},N^{*})$ non-trivially in the event that either of its endpoints belong to $\partial^{s}R^{\prime}_{k_{0}+1}(\lambda,Q_{j},N^{*})$. However, from (3.19), and since we chose $N^{*}$ so large that $R^{\prime}_{k_{0}+1}(\lambda,Q_{j},N^{*})\subset B_{\frac{C^{**}}{4}}$ (see (3.7)), it must be the case that $f_{\lambda,k_{0},0}\left(I_{0}^{\prime}\right)$ indeed connects the stable boundaries of $R^{\prime}_{k_{0}+1}(\lambda,Q_{j},N^{*})$. In particular, by Lemma 13, if $I_{0}^{\prime}$ has endpoints $P_{1}$ and $P_{2}$ with $P_{1}<P_{2}$, then we have $f_{\lambda,k_{0},0}(P_{1})\in\partial^{s}_{-}R^{\prime}_{k_{0}+1}(\lambda,Q_{j},N^{*})$ and $f_{\lambda,k_{0},0}(P_{2})\in\partial^{s}_{+}R^{\prime}_{k_{0}+1}(\lambda,Q_{j},N^{*})$.
Suppose that we have constructed all of the gaps up to order $k\geq k_{0}$, and that they are all disjoint segments of $\hat{L}_{\lambda}$ of the form $\text{int}I_{m}$ where $I_{m}\subset I_{0}$ for some $m\leq k$ such that:
@@
The gaps of order $k+1$ will come from taking the image of such components that intersects $\bigcup_{i=1}^{4}O_{r_{0}}(Q_{i})$ non-trivially, under the map $f^{-1}_{\lambda,k+1,0}$. A similar application of Remark 5 and Lemma 13 as in the base case shows that any gap of order $k+1$ will be of the form above with index $m$ replaced with $k+1$.
@@
$C\left(\lambda,N^{*}\right)\subset C\left(\lambda,O_{r(\rho)}\right).$ (3.20)
3.6.2 Thickness Estimate
@@
$\geq C_{4}^{-1}\frac{\text{dist}_{f_{\lambda,k,0}\left(\hat{L}_{\lambda}\right)}\left(f_{\lambda,k,0}(P),f_{\lambda,k,0}\left(Q\right)\right)}{\text{dist}_{f_{\lambda,k,0}\left(\hat{L}_{\lambda}\right)}\left(f_{\lambda,k,0}(P),f_{\lambda,k,0}\left(P^{\prime}\right)\right)}\geq C_{4}^{-1}\frac{\frac{C^{**}}{4}}{C_{1}C^{*}\left(\mu_{1}-\delta\right)^{-N}}-C_{4}^{-1}$
@@
From (3.20), this then implies (3.11) holds, and hence (3.1) holds, for all $\lambda\in[0,\lambda_{0}]$.
Since $L_{\lambda}\cap\Lambda^{\text{bnd}}_{\lambda}$ is affine equivalent to $\sigma_{\lambda,\alpha}$ and $M$ was arbitrary, this proves Theorem 1.
@@ Acknowledgements
I would like to thank Anton Gorodetski for his guidance and support. This project was supported by NSF grant DMS-2247966 (PI: A. Gorodetski).

### v2

@@ 4.6.2 Rectangles of Controlled Size
We will now use this base set of rectangles to construct dynamical rectangles of small controlled size. We will need rectangles of smaller scale to ensure that the sub-Cantor set of $\mathcal{C}(\lambda,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{O}_{r(\rho)}})$, that we construct later on, will be thick and hence have Hausdorff dimension larger than $\frac{\log 2}{\log\left(2+\frac{1}{M}\right)}$.
@@
    $\text{length}(\kappa)\leq C_{1}(\mu_{1}-\delta)^{-N}C^{*}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}.}$
@@
    $\text{length}(\kappa)\leq C_{1}(\mu_{1}-\delta)^{-N}\frac{C^{*}}{2}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}.}$
@@
    The non-stable boundaries $\partial^{ns}R^{\prime}_{m}(\lambda,Q_{i},N)$ of $R^{\prime}_{m}(\lambda,Q_{i},N)$ are line segments with endpoints in $\partial\tilde{K}(Q_{i})$. That is, $R^{\prime}_{m}(\lambda,Q_{i},N)$ has vertices belonging to $\partial\tilde{K}(Q_{i})$, and does not intersect ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\partial}\tilde{K}(Q_{i})$ otherwise.
@@
Notice from (4.12) and Lemma 6, provided $\lambda$ is small, the image of the cone $\tilde{K}(Q_{i})$, under the map $f_{\lambda,0,m}$, overflows the box $\mathbb{B}_{j}$ (see Figure 10). Without loss of generality, let us assume that this holds for all $\lambda\in[0,\lambda_{4}].$
@@ 4.7 Gap Estimates and Thickness
In order to establish (4.9), we will show that for each $\lambda\in[0,\lambda_{4}]$, there is a Cantor set $\mathcal{C}\left(\lambda,N^{*}\right)$ contained in $\mathcal{C}\left(\lambda,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{O}_{r(\rho)}}\right)$ that has Hausdorff dimension larger than $\frac{\log 2}{\log\left(2+\frac{1}{M}\right)}$.
4.7.1 Thickness of Cantor Sets
A fundamental tool we will use to estimate the Hausdorff dimension of this set is the notion of thickness for Cantor sets. Given a Cantor set $\mathcal{C}\subset\mathbb{R}$ with convex hull $I$, we say that $U\subset I\setminus\mathcal{C}$ is a gap of $\mathcal{C}$ if it is a connected component of $I\setminus\mathcal{C}$. Any enumeration $\mathcal{U}=\{U_{n}\}$ of the gaps of $\mathcal{C}$ is called a presentation of $\mathcal{C}$. Given a boundary point $u\in\mathcal{C}$ of some gap $U_{k}$, we call the connected component $B$ contained in $I\setminus\left(U_{1}\cup\dots\cup U_{k}\right)$ that contains $u$ a bridge of $U_{k}$. We define
$\tau(\mathcal{C},\mathcal{U},u):=\frac{|B|}{|U_{k}|},$
and set the thickness of $\mathcal{C}$ to be
(see Chapter 4 of [PT]). It is clear that these notions carry over to a Cantor set that belongs to a $C^{1}$ curve $\gamma\subset\mathbb{T}^{2}$ where the lengths of gaps and bridges are measured using length relative to the curve $\gamma$.
4.7.2 Ordering of $\hat{L}_{\lambda}$
@@ 4.7.3 Gap Constructions
We will construct a Cantor set $\mathcal{C}\left(\lambda,N^{*}\right)$ that is contained in $\mathcal{C}\left({\lambda},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{O}_{r(\rho)}}\right)$ and show that it has thickness bigger than $M$:
@@
Let $J$ be such a component, and suppose it is contained in $R^{\prime}_{k_{0}+1}(\lambda,Q_{j},N^{*})$ for some $j=1,\dots,4$, and $J\cap O_{r_{0}}(Q_{j})\neq\emptyset$. Then, $J$ is the image of some segment $I_{0}^{\prime}\subset I_{0}$ under the map $f_{\lambda,k_{0},0}$. We will say that ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{int}\left(I^{\prime}_{0}\right)}$ is a gap of order $k_{0}$, where the interior is taken with respect to the subspace topology on $\hat{L}_{\lambda}$.
@@
$f_{\lambda,k_{0},0}\left(I_{0}^{\prime}\right)$ is a curve tangent to $\mathcal{K}^{u}$, by Proposition 2 and our discussion after Lemma 16, as $f_{\lambda,k_{0},0}\left(I_{0}^{\prime}\right)\cap O_{r_{0}}(Q_{j})\neq\emptyset$, we must have that $f_{\lambda,k_{0},0}\left(I_{0}^{\prime}\right)$ is fully contained in $R^{\prime}_{k_{0}+1}(\lambda,Q_{j},N^{*})$ and only intersects $\partial R^{\prime}_{k_{0}+1}(\lambda,Q_{j},N^{*})$ non-trivially in the event that either of its endpoints belong to $\partial^{s}R^{\prime}_{k_{0}+1}(\lambda,Q_{j},N^{*})$. However, from (4.19), and since we chose $N^{*}$ so large that $R^{\prime}_{k_{0}+1}(\lambda,Q_{j},N^{*})\subset B_{\frac{C^{**}}{4}}$ (see (4.7)), it must be the case that $f_{\lambda,k_{0},0}\left(I_{0}^{\prime}\right)$ indeed connects the stable boundaries of $R^{\prime}_{k_{0}+1}(\lambda,Q_{j},N^{*})$. In particular, by Lemma 17, if $I_{0}^{\prime}$ has endpoints $P_{1}$ and $P_{2}$ with $P_{1}<P_{2}$, then we have $f_{\lambda,k_{0},0}(P_{1})\in\partial^{s}_{-}R^{\prime}_{k_{0}+1}(\lambda,Q_{j},N^{*})$ and $f_{\lambda,k_{0},0}(P_{2})\in\partial^{s}_{+}R^{\prime}_{k_{0}+1}(\lambda,Q_{j},N^{*})$.
Suppose that we have constructed all of the gaps up to order $k\geq k_{0}$, and that they are all disjoint segments of $\hat{L}_{\lambda}$ of the form ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{int}\left(I_{m}\right)}$ where $I_{m}\subset I_{0}$ for some $m\leq k$ such that:
@@
The gaps of order $k+1$ will come from taking the image of such components that intersects $\bigcup_{i=1}^{4}O_{r_{0}}(Q_{i})$ non-trivially, under the map $f^{-1}_{\lambda,k+1,0}$. A similar application of Proposition 2 and Lemma 17 as in the base case shows that any gap of order $k+1$ will be of the form above with index $m$ replaced with $k+1$.
@@
$C\left(\lambda,N^{*}\right)\subset C\left(\lambda,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{O}_{r(\rho)}}\right).$ (4.20)
4.8 Concluding the Proof of Theorem 4
With the construction of the set $\mathcal{C}\left(\lambda,N^{*}\right)$, we are now ready to complete the proof of Theorem 4.
Proof of Theorem 4..
@@ 4.8 Concluding the Proof of Theorem 4
$\geq C_{4}^{-1}\frac{\text{dist}_{f_{\lambda,k,0}\left(\hat{L}_{\lambda}\right)}\left(f_{\lambda,k,0}(P),f_{\lambda,k,0}\left(Q\right)\right)}{\text{dist}_{f_{\lambda,k,0}\left(\hat{L}_{\lambda}\right)}\left(f_{\lambda,k,0}(P),f_{\lambda,k,0}\left(P^{\prime}\right)\right)}\geq C_{4}^{-1}\frac{\frac{C^{**}}{4}}{C_{1}C^{*}\left(\mu_{1}-\delta\right)^{-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}N^{*}}}}-C_{4}^{-1}$
@@
From (4.20), this then implies (4.9) holds, for all $\lambda\in[0,\lambda_{0}]$. ∎
5 Concluding the Proofs
Proof of Theorem 3..
Fix a bounded sequence $\overline{a}=(a_{n})$, $a_{n}\in\mathbb{N}$, and denote $\mathcal{A}:=\{a_{n}\}$. Let $\epsilon>0$ and suppose that $M>0$ is so large that
$\frac{\log 2}{\log\left(2+\frac{1}{M}\right)}>1-\epsilon.$
By Theorem 4, there is a $\rho>0$ and $\lambda_{0}\in\left(0,\min_{a\in\mathcal{A}}\left\{\tilde{\lambda}(\rho,a)\right\}\right)$ such that $\mathcal{C}\left(\lambda,\tilde{O}_{r(\rho)}\right)$ has Hausdorff dimension larger than $\frac{\log 2}{\log\left(2+\frac{1}{M}\right)}$.
Set $\lambda\in(0,\lambda_{0})$. Then, since $\lambda_{0}<\min_{a\in\mathcal{A}}\left\{\tilde{\lambda}(\rho,a)\right\}$, from Lemma 5, we know that $\Lambda_{\lambda}^{\text{surv}}\left(\overline{a}\right)\cap L_{\lambda}$ is the image of $\mathcal{C}\left(\lambda,\tilde{O}_{r(\rho)}\right)$ under the map $F_{\lambda}$. Since $F$ is a local diffeomorphism, it follows that
$\text{dim}_{H}\left(\Lambda_{\lambda}^{\text{surv}}\left(\overline{a}\right)\cap L_{\lambda}\right)=\text{dim}_{H}\left(\mathcal{C}\left(\lambda,\tilde{O}_{r(\rho)}\right)\right)\geq 1-\epsilon.$
Proof of Theorem 1..
Let $\alpha=[a_{1},a_{2},\dots]\in(0,1)$ be an irrational of bounded type. Set $\overline{a}=(a_{n}).$ Let $\epsilon>0$. By Theorem 3, there is a $\rho>0$ and $\lambda_{0}>0$ such that for any $\lambda\in[0,\lambda_{0}]$, we have
$\text{dim}_{H}\left(\Lambda_{\lambda}^{\text{surv}}\left(\overline{a}\right)\cap L_{\lambda}\right)\geq 1-\epsilon.$
Let $\lambda\in[0,\lambda_{0}]$. From Theorem 2, and since $\Lambda^{\text{surv}}_{\lambda}\left(\overline{a},O_{\rho}\right)\subset\Lambda_{\lambda}^{\text{bnd}}\left(\overline{a}\right)$, we have
$\text{dim}_{H}\left(\sigma_{\lambda,\alpha}\right)=\text{dim}_{H}\left(L_{\lambda}\cap\Lambda_{\lambda}^{\text{bnd}}\left(\overline{a}\right)\right)\geq\text{dim}_{H}\left(\Lambda_{\lambda}^{\text{surv}}\left(\overline{a}\right)\cap L_{\lambda}\right)\geq 1-\epsilon.$
Remark 6.
We expect that Theorem 1 should hold for all irrational $\alpha\in(0,1)$. We would like to point out exactly where the assumption that $\alpha$ is of bounded type is used in our proofs. It is used in Proposition B.1, which is the main ingredient of Lemma 4, and this is what later allows us to control the distance of each piece of stable manifold $W^{s}_{\lambda}(Q_{0},m)$ from each $Q_{i}$ in the construction of the rectangles in Subsection 4.6. The assumption also implies that the collection of maps $\left\{\hat{T}_{\lambda,a}:\lambda\in[0,\lambda_{0}],\ a\in\mathcal{A}\right\}$ from Section 4 have uniformly bounded $C^{2}$ norms, which is required for the bounded distortion estimate (see Lemma 14). It also allows us to control the $C^{2}$-closeness of $\hat{T}_{\lambda,a}$ and $\hat{T}_{0,a}$ in a uniform way, which is used in many arguments such as Theorem 5 and Corollary 1.
Appendix A Contraction Mappings
Proposition A.1.
$d\left(f_{n}(x),f_{n}(y)\right)\leq\mu d(x,y),$
for all $x,y\in K$. Then, the following hold:
  • (1)
    There is a $x^{*}\in K$ such that
    for any $x\in K$.
  • (2)
    $d\left(h_{n}(x),h_{n}(y)\right)\leq\mu^{\prime}d(x,y)\ \text{and}\ \sup_{n}d_{C^{0}(K)}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{n}},h_{n})<\epsilon^{\prime},$
    for some $\mu^{\prime}\in(0,1)$ and for all $x,y\in K$, then
    $d\left(x^{*},y^{*}\right)<\epsilon,$
    where $y^{*}$ is the element such that
    for each $x\in X$.
Proof.
  • (1)
    The proof here is analogous to that of the Banach Fixed Point Theorem. For a proof of this exact result, see Theorem 3 in [LY].
  • (2)
    for any $x\in K$. We can deduce that
Appendix B Continued Fractions
$\frac{p_{n}(\alpha)}{q_{n}(\alpha)}=[a_{1},a_{2},\dots,a_{n}].$
$p_{k+1}(\alpha)=a_{k+1}p_{k}(\alpha)+p_{k-1}(\alpha),\ p_{1}(\alpha)=p_{0}(\alpha)=1$ (B.1)
$q_{k+1}(\alpha)=a_{k+1}q_{k}(\alpha)+q_{k-1}(\alpha),\ q_{1}(\alpha)=a_{1},\ q_{0}(\alpha)=1$ (B.2)
and that
$\left|q_{k}(\alpha)\alpha-p_{k}(\alpha)\right|<\frac{1}{q_{k}(\alpha)}.$ (B.3)
and by (B.3), both of these quantities are no greater than $\frac{1}{q_{k-1}(\alpha)}+\frac{1}{q_{k}(\alpha)}$. As a corollary, we obtain the following result which will be used later on.
Proposition B.1.
$\min\limits_{1\leq m\leq n}\left|m\alpha-y\right|<\epsilon$
for any $y\in S^{1}$.
Proof.
Let $\epsilon>0.$ Set $Y:=\{\alpha=[a_{1},a_{2},\dots]:a_{i}\in\mathcal{A}\}\subset[0,1]\setminus\mathbb{Q}$ and let $a^{\prime}$ be the smallest element of $\mathcal{A}$. Then using the recurrence relation (B.2) and an induction argument, we find that for each $k\geq 1$, there is an $M_{k}>0$ such that
Choose $k$ large enough so that
@@ Acknowledgements
I would like to thank Anton Gorodetski for his guidance and support, and Jake Fillman for looking at a preliminary draft of this text. I would also like to thank both anonymous referees for their helpful suggestions and input. This project was supported by NSF grant DMS-2247966 (PI: A. Gorodetski).